# Performance Rating Equilibrium

### v2

@@ Abstract
In this note, I introduce a novel performance rating system called Performance Rating Equilibrium (PRE). A PRE is a vector of hypothetical ratings for each player, such that if these ratings were each player’s initial rating at the start of a tournament, scoring the same points against the same opponents would leave each player’s initial rating unchanged. In other words, all players’ initial ratings perfectly predict their actual scores in the tournament. This property, however, does not hold for the well-known Tournament Performance Rating. PRE is defined as a fixed point of a multidimensional rating function. I show that such a fixed point, and hence a PRE, exists under mild conditions. I provide an implementation of PRE along with several empirical applications. PREs have broad applicability, from sports competitions to the evaluation of large language models. JEL: Z20, D63
@@ 1 Introduction
Rating and ranking systems are widely studied across various fields, including economics, social choice, sports, and computer science, to assess individual performance or rank objects. One prominent approach, the Elo rating system, originally developed for chess by Elo (1978), has been applied in diverse domains, including association football (FIFA, 2018; Csató, 2021), tennis (Williams et al., 2021), college rankings (Avery et al., 2012), and large language model evaluation (Zheng et al., 2024). [^2] Alternative systems, such as the Glicko rating system (Glickman, 1995), may also be used for ratings in these areas.
In a chess tournament, players’ scores against opponents with specific ratings are used to calculate their Tournament Performance Rating (TPR). The TPR is the hypothetical rating $R$ that would remain unchanged after the tournament. However, because TPR is calculated directly from opponents’ pre-tournament ratings, it does not take into account their actual performances in the event.
@@
Here, I introduce the concept of Performance Rating Equilibrium (PRE), represented by a vector $r^{*}$ of hypothetical ratings. A PRE is defined as a fixed point of a multidimensional rating function: if each player $i$ began the tournament with an initial rating $r^{*}_{i}$, then achieving the same points against the same opponents would leave each player’s initial rating $r^{*}_{i}$ unchanged. In other words, these initial ratings perfectly predict each player’s final score, and I refer to each player’s rating at a PRE as a perfect performance rating (PPR).
@@
Landau (1895, 1914) introduced one of the earliest score-based ranking systems inspired by round-robin chess tournaments. Iterative methods for ranking, whether of players or political candidates, date back to at least Wei (1952); Kendall (1955), who extended an approach by Wei (1952). For some selected further developments, see Keener (1993); Laslier (1997); Herings et al. (2005); Brozos-Vázquez et al. (2008); González-Díaz et al. (2014); Csató (2015, 2017), as well as the influential PageRank algorithm introduced by Brin and Page (1998). Herings et al. (2005) propose a new method for ranking the nodes in directed graphs, while González-Díaz et al. (2014) provide an axiomatic analysis of several ranking methods. [^3] Rubinstein (1980) provides an axiomatization of (non-iterative) score-based ranking. In addition, Csató (2017) explores Swiss-system team tournaments, proposing various score-based ranking methods and applying them to the European Chess Team Championships. [^4] For a broader, growing literature on fairness in sports, see, e.g., Brams and Ismail (2018); Brams et al. (2018); Csató (2021); Lambers and Spieksma (2021); Palacios-Huerta (2023); Devriesere et al. (2024), and the references therein.
@@
The early pioneering work by Kendall (1955) focused on recursive rankings that incorporate both the players’ current performances (measured by scores) and the performances of their opponents in the tournament. To my knowledge, the study by Brozos-Vázquez et al. (2008) is the first to use performance ratings, called “Recursive Performance,” that account for the opponents’ current performance ratings rather than only current scores in their ranking method. There are three main differences between Recursive Performance and the current study. First, Brozos-Vázquez et al.’s main focus is on ranking participants, whereas the focus here is on finding ratings that perfectly predict the actual performances of players. Second, although Recursive Performance offers an improvement over the standard TPR, it, like TPR, does not generally yield a fixed point of the rating function, so these concepts are not in equilibrium as defined here. [^5] The equilibrium Lastly, Recursive Performance ratings are calculated using average opponent ratings with a corrective factor based on score. Although this approach approximates actual performance ratings reasonably well when rating gaps are small, it is well known that it can lead to paradoxical outcomes, such as a player’s performance rating decreasing after winning against a low-rated opponent.
@@ 2 Setup
Let $T=(N,r,k,m,f)$ denote a tournament, where $N=\{1,2,\dots,n\}$ is the set of players. The non-negative vectors $r\in\mathbb{R}^{n}_{+}$, $k\in\mathbb{N}^{n}_{+}$, and $m\in\mathbb{R}^{n}_{+}$ denote the initial ratings, the number of games played by the players, and their final scores, respectively, with $0\leq m_{i}\leq k_{i}$ for each $i\in N$. Here, $m_{i}$ denotes $i$’s (final) score in $k_{i}$ games. The function $f:\mathbb{R}^{n}_{+}\rightarrow\mathbb{R}^{n}_{+}$ is called a rating function. A vector of ratings $x^{*}\in\mathbb{R}^{n}_{+}$ is called a Performance Rating Equilibrium (PRE) if $f(x^{*})=x^{*}$. [^6] Here, the term ‘equilibrium’ is used in the sense of a steady state; it is an interactive but a non-choice-based concept.
Let $[f(\cdot)]_{a}^{b}=\min(\max(f(\cdot),a),b)$ denote a function bounded within the interval $[a,b]$. Each player $i$ plays $k_{i}$ games. Denote $i$’s opponents in these games by $i_{1},i_{2},\dots,i_{k_{i}}$. For some $j$ and $j^{\prime}$, it may be that $i_{j}=i_{j^{\prime}}$, meaning that a player may play more than once against the same opponent. Let
$c=\max_{i\in N}\sum_{j=1}^{k_{i}}r_{i_{j}}$
denote the highest possible value for the sum of ratings of a player’s opponents. The role of $c$ is to provide a bound for the rating function. While the main result holds for any upper bound $c>0$ and any chosen lower bound, the definition of $c$ is arguably a reasonable bound for a rating function based on initial ratings $r$ because, even if a player wins all games, the player’s new rating should not strictly exceed the sum of the ratings of the player’s opponents.
Fix a tournament $T$. Note that Tournament Performance Rating (TPR) is undefined for zero or perfect scores, as Elo himself noted (Elo, 1978, p. 12). Ismail (2023) introduces Complete Performance Rating (CPR), which is defined for all scores, including when $m_{i}=0$ or $m_{i}=k_{i}$.
For $x\in\mathbb{R}^{n}_{+}$ and for every player $i$ with $0<m_{i}<k_{i}$, let $y_{i}$ be the value that solves this equation
$m_{i}=\sum_{j=1}^{k_{i}}\frac{1}{1+10^{\frac{x_{i_{j}}-y_{i}}{400}}}.$ (1)
Define $\text{TPR}_{i}(x)=[y_{i}]_{0}^{c}$. For $m_{i}=0$ or $m_{i}=k_{i}$, define $\text{TPR}_{i}(x)=\left[\text{CPR}_{i}(x)\right]_{0}^{c}$.
The following theorem establishes the existence of PREs in tournaments.
Main Theorem: Let $T=(N,r,k,m,\text{TPR})$ be a tournament. Then, there exists a PRE in $T$.

### v3

@@ Abstract
In this note, we introduce a novel performance rating system called Performance Rating Equilibrium (PRE). A PRE is a vector of ratings for each player, such that if these ratings were each player’s initial rating at the start of a tournament, scoring the same points against the same opponents would leave each player’s initial rating unchanged. In other words, all players’ initial ratings perfectly predict their actual scores in the tournament. This property, however, does not hold for the well-known Tournament Performance Rating. PRE is defined as a fixed point of a multidimensional rating function. We show that such a fixed point, and hence a PRE, exists under mild conditions. We provide an implementation of PRE along with several empirical applications. PREs have broad applicability, from sports competitions to the evaluation of large language models. JEL: Z20, D71, D63
@@ 1 Introduction
Rating and ranking systems are widely studied across various fields, including (political) economics, social choice, sports, and computer science, to assess individual performance or rank objects. One prominent approach, the Elo rating system, originally developed for chess by Elo (1978), has been applied in diverse domains, including association football (FIFA, 2018; Csató, 2021), tennis (Williams et al., 2021), college rankings (Avery et al., 2013), dating applications (Kosoff, 2016), and large language model evaluation (Zheng et al., 2024). [^1] Alternative systems, such as the Glicko rating system (Glickman, 1995), may also be used for ratings in these areas.
In the Elo rating system, players are assigned a numerical rating, which is updated based on their performance over a given period. A player gains Elo rating if his or her score against opponents is better than the expected score predicted by the Elo ratings. Since Elo rating is cumulative, based on all games that a player has played, it does not necessarily reflect a player’s recent performance. The Tournament Performance Rating (TPR) is the standard method for calculating a player’s performance in a subset of games, such as a specific tournament. [^2] In the case of LLM evaluation, such as in the Chatbot Arena, the TPR of an LLM can be calculated based on a single-day performance or a subject-specific performance, such as mathematics. A player’s TPR is the hypothetical rating $R$ that would remain unchanged based on the player’s performance in a tournament. However, because TPR is calculated directly from opponents’ pre-tournament ratings, it does not take into account their actual performances in the event.
@@
Here, we introduce the concept of Performance Rating Equilibrium (PRE), represented by a vector $x^{*}$ of hypothetical ratings. A PRE is defined as a fixed point of a multidimensional rating function: if each player $i$ began the tournament with an initial rating $x^{*}_{i}$, then achieving the same points against the same opponents would leave each player’s initial rating $x^{*}_{i}$ unchanged. In other words, these initial ratings perfectly predict each player’s final score, and we refer to each player’s rating at a PRE as a perfect performance rating (PPR). [^3] The term ‘equilibrium’ here refers to an equilibrium point or steady state of a dynamical system. This concept is not choice-based, though it is interactive, as each PPR in a PRE depends on the other PPRs.
@@
Landau (1895, 1914) introduced one of the earliest score-based ranking systems inspired by round-robin chess tournaments. Iterative methods for ranking, whether of players or political candidates, date back to at least Wei (1952); Kendall (1955), who extended an approach by Wei (1952). For some selected further developments, see Keener (1993); Laslier (1997); Herings et al. (2005); Brozos-Vázquez et al. (2008); González-Díaz et al. (2014); Csató (2015, 2017), as well as the influential PageRank algorithm introduced by Brin and Page (1998). Herings et al. (2005) propose a new method for ranking the nodes in directed graphs, while González-Díaz et al. (2014) provide an axiomatic analysis of several ranking methods, including generalised row sum methods by Chebotarev (1994). [^4] Rubinstein (1980) provides an axiomatization of (non-iterative) score-based ranking. In addition, Csató (2017) explores Swiss-system team tournaments, proposing various score-based ranking methods essentially based on generalised row sum methods and applying them to the European Chess Team Championships. [^5] For a broader, growing literature on fairness in sports, see, e.g., Brams and Ismail (2018); Brams et al. (2018); Csató (2021); Lambers and Spieksma (2021); Palacios-Huerta (2023); Krumer and Moreno-Ternero (2023); Csató et al. (2024); Devriesere et al. (2024), and the references therein.
@@
The early pioneering work by Kendall (1955) focused on recursive rankings that incorporate both the players’ current performances (measured by scores) and the performances of their opponents in the tournament. To my knowledge, the study by Brozos-Vázquez et al. (2008) is the first to use performance ratings, called “Recursive Performance,” that account for the opponents’ current performance ratings rather than only current scores in their ranking method. There are three main differences between Recursive Performance and the current study. First, Brozos-Vázquez et al.’s main focus is on ranking participants, whereas the focus here is on finding ratings that perfectly predict the actual performances of players. Second, although Recursive Performance offers an improvement over the standard TPR, it, like TPR, does not generally yield a fixed point of the rating function, so these concepts are not in equilibrium as defined here. Lastly, Recursive Performance ratings are calculated using average opponent ratings with a corrective factor based on score. Although this approach approximates actual performance ratings reasonably well when rating gaps are small, it is well known that it can lead to paradoxical outcomes, such as a player’s performance rating decreasing after winning against a low-rated opponent.
@@ 2 Setup
Let $T=(N,Y,W,f,\psi)$ denote a tournament, where:
  1. 1.
    $N=\{1,2,\dots,n\}$ is the set of players.
  2. 2.
    $W=(w_{ij})_{i,j\in N}$ is the symmetric schedule matrix, where $w_{ij}=w_{ji}\geq 0$ denotes the number (or weight) of games between players $i$ and $j$, and $w_{ii}=0$.
  3. 3.
    $Y=(y_{ij})_{i,j\in N}$ with $0\leq y_{ij}\leq w_{ij}$ denotes realized points earned by $i$ against $j$, with
[…]
  4. 4.
    $k\in\mathbb{R}_{\geq 0}^{n}$ is the vector of games played by each player, derived from $W$:
    $k_{i}:=\sum_{j\neq i}w_{ij}.$
  5. 5.
    $m\in\mathbb{R}_{\geq 0}^{n}$ is the vector of scores, derived from $Y$:
    $m_{i}:=\sum_{j\neq i}y_{ij}.$
    The total score is
    $M:=\sum_{i=1}^{n}m_{i}\;=\;\sum_{i<j}w_{ij}.$
    We assume $0<m_{i}<k_{i}$ for all $i$ with $k_{i}>0$.
  6. 6.
    $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the performance rating function associated with $T$. For any hypothetical initial ratings $x\in\mathbb{R}^{n}$, the vector $f(x)$ returns the performance ratings implied by $T$.
  7. 7.
    Finally, $\psi:\mathbb{R}\to(0,1)$ (defined below) is a monotone antisymmetric function that gives the win probability (or the expected score share) in a single game as a function of the rating difference.
Definition 1 (Win Probability).
The win probability function $\psi:\mathbb{R}\to(0,1)$ models the expected score share of a player in a single game against an opponent, depending on the difference in their ratings. We assume:
  1. (a)
    $\psi$ is $C^{1}$ and strictly increasing: $\psi^{\prime}(t)>0$ for all $t$;
  2. (b)
    $\psi(-t)=1-\psi(t)$ for all $t$ (antisymmetry);
  3. (c)
    $\lim_{t\to+\infty}\psi(t)=1$ and $\lim_{t\to-\infty}\psi(t)=0$.
Definition 2 (Performance Rating Equilibrium).
Given a tournament $T$, a vector $x^{*}\in\mathbb{R}^{n}$ is a Performance Rating Equilibrium (PRE) if it is a fixed point of the performance rating function $f$:
$f(x^{*})=x^{*}.$
We call $x_{i}^{*}$ the perfect performance rating (PPR) of player $i$ at the PRE.
Definition 3 (Expected Score).
Given $x\in\mathbb{R}^{n}$, write $\Delta_{ij}(x):=(x_{i}-x_{j})/\beta$, where $\beta>0$, and define the model-implied expected (total) score of player $i\in N$ in tournament $T$ by
$E_{i}(x):=\sum_{j\neq i}w_{ij}\,\psi\big(\Delta_{ij}(x)\big).$
Let $E(x)=(E_{i}(x))_{i\in N}$.
Remark 1.
Antisymmetry and differentiability imply $\psi^{\prime}(-t)=\psi^{\prime}(t)$, so $\psi^{\prime}$ is even. Moreover, only rating differences matter: for any $c\in\mathbb{R}$, $E(x+c\mathbf{1})=E(x)$, where $\mathbf{1}$ the vector of all ones.
Lemma 1 (PRE equivalence).
There exists a choice of $f$, the canonical $\psi$–performance function, such that
$f(x)=x\quad\Longleftrightarrow\quad E(x)=m.$
Therefore, the PRE condition is exactly $E(x^{*})=m$.